# When Benchmarks are Targets: Revealing the Sensitivity of Large Language Model Leaderboards

### v1

@@ Abstract
Large Language Model (LLM) leaderboards based on benchmark rankings are regularly used to guide practitioners in model selection. Often, the published leaderboard rankings are taken at face value — we show this is a (potentially costly) mistake. Under existing leaderboards, the relative performance of LLMs is highly sensitive to (often minute) details. We show that for popular multiple choice question benchmarks (e.g. MMLU) minor perturbations to the benchmark, such as changing the order of choices or the method of answer selection, result in changes in rankings up to 8 positions. We explain this phenomenon by conducting systematic experiments over three broad categories of benchmark perturbations and identifying the sources of this behavior. Our analysis results in several best-practice recommendations, including the advantage of a hybrid scoring method for answer selection. Our study highlights the dangers of relying on simple benchmark evaluations and charts the path for more robust evaluation schemes on the existing benchmarks.
@@
Norah Alzahrani††thanks: Core contributor, Hisham Abdullah Alyahya††footnotemark: , Yazeed Alnumay, Sultan Alrashed, Shaykhah Alsubaie, Yusef Almushaykeh, Faisal Mirza, Nouf Alotaibi Nora Altwairesh, Areeb Alowisheq, M Saiful Bari, Haidar Khan††footnotemark: ††thanks: Corresponding author: haidark@sdaia.gov.sa National Center for AI, Saudi Arabia
@@ 1 Introduction
The advent of transformer-based Large Language Models (LLMs) (OpenAI, 2023; Deepmind, 2023; Anthropic, 2023; Anil et al., 2023; Touvron et al., 2023) has led to a generational leap in generative models, enabling interaction with computing devices through natural language. This advancement encompasses improvements that have rendered many earlier benchmarks and leaderboards obsolete (Laskar et al., 2023; Shen et al., 2023), leading to the compilation of more challenging and comprehensive tests. However, the current generation of leaderboards still do not satisfy many of the requirements of researchers and practitioners looking to build on LLMs Ethayarajh and Jurafsky (2021); Dehghani et al. (2021). Since LLMs are extremely expensive to both train and inference, selecting the LLM (or LLM training recipe) is often the most costly decision for the entire project. Stable leaderboards are critical to making the right decision.
@@
Leaderboards based on multiple choice questions (MCQ) for evaluation (Wang et al., 2018, 2019; Nie et al., 2019; Zhong et al., 2023; Hendrycks et al., 2020) present both convenience and significant limitations (Pezeshkpour and Hruschka, 2023; Zheng et al., 2023). While MCQs offer a seemingly straightforward, automated, and quantifiable means to assess certain aspects of model ability (e.g. knowledge), they fall short as a stable means to measure performance. Figure 1 demonstrates the instability of the leaderboard ranking of one popular benchmark, Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020), under small perturbations.
@@
Moreover, the reliance on MCQs raises concerns about the models being overfit to these benchmarks, potentially excelling in structured tests while lacking real-world applicability. This discrepancy highlights the need for more holistic and diverse evaluation methods that transcend the simplicity of MCQs Liang et al. (2023). It also prompts critical reflection on how these models might inadvertently be trained to cheat the tests – achieving high scores through pattern recognition and optimization for specific question formats rather than genuine language comprehension or knowledge. As LLMs continue to evolve, it is imperative to develop evaluation frameworks that can more accurately assess their abilities in a way that mirrors the complexity of real-world use.
@@
Despite being widely used, benchmarking with MCQs has turned out to be anything but simple. It requires the full synchronization of evaluation frameworks and results often vary wildly due to nuanced differences. For example, minor changes in prompting and scoring can produce invalid results for particular LLMs111https://huggingface.co/blog/evaluating-mmlu-leaderboard. Two recent studies demonstrate that LLMs are susceptible to the ordering of answer choices and bias towards specific tokens/symbols (Zheng et al., 2023; Pezeshkpour and Hruschka, 2023). In this work, we observe how minor perturbations to MCQ can disrupt model rankings on leaderboards based on MCQ benchmarks. We also take additional steps to precisely identify the limitations of LLMs on this measurement approach.
@@
Demonstrate that some categories of modifications do not affect the benchmark rankings.
@@ 2 LLM Evaluation with MCQs
Mechanically, testing LLMs wth MCQs is accomplished by presenting the question along with the answer choices to the model and selecting the choice deemed most probable by the model. Although this setup appears straightforward, LLMs react in unpredictable ways to formatting and other minor changes to the questions or the answers. LLM performance on a MCQ test can change with the introduction of an extra space (e.g. between the question and answer) or adding an additional instructional phrase (e.g. "Choices:"). In addition to this brittleness, Pezeshkpour and Hruschka (2023) found changes to the order in which answer choices are presented to GPT4 and instructGPT can change the model’s prediction.
@@
Our main aim is to quantify how these small perturbations/variations change the rankings of a set of models on a particular benchmark. As MCQ benchmarks based leaderboards are often used to compare models and guide model selection, we investigate the robustness of benchmarks for this purpose. Figure 1 demonstrates how existing benchmarks exhibit significant undesirable shifts in rankings under small perturbations.
@@ 3 Methods
In this section, we describe and justify the pertubrations we apply in each of the categories. We note that some MCQ tests changes, like modifying the order of answer choices can change perfomance even for humans but the effect is typically not pronounced (Lions et al., 2021). In general, the modifications we make are designed to be small perturbations to the MCQ and prompts that should not affect performance. The exception to this are some of the in-context knowledge manipulations described in Section 3.3, which are designed to drastically improve or degrade performance.
@@ 3.1 Answer choice format and ordering
In light of earlier findings related to selection bias Zheng et al. (2023); Pezeshkpour and Hruschka (2023), we investigate the effects of changes to the presented order of answer choices and changes to the symbols associated with answer choices.
@@
Our first study aims to uncover how dependent MCQ benchmark performance and rankings are on the original ordering of the answer choices. We apply two simple schemes to randomly change the order of answer choices presented to the model: (i) swapping choices using a fix set of swaps for all questions and (ii) randomly assigning new positions to each choice while ensuring each choice is moved to a different position.
@@
In this setting, the correct answer choice is set to a fixed position across the entire test to measure bias toward predicting answers at particular positions. For 0-shot, we simply set the correct answer choice to each of the positions in turn.
In the few-shot case, we examine the influence of biasing the correct answers in the examples to the model’s inherent bias to particular positions. For each question, we fix the correct answer of the examples to each position in turn. We then modify the test question in two ways: (i) unchanged answer choicess and (ii) correct choice fixed to same position as examples. This setup is shown in Figure 2.
@@
The symbols used for the answer choices (e.g. A, B, C, D) also play a role in model bias Zheng et al. (2023), thus we experiment with replacing the symbols with alternative and less common tokens. The goal of this is to decouple the bias to particular positions from the bias to symbols or the relative ordering in natural symbols. We replace [’A’,’B’,’C’,’D’] with the following two sets of symbols:
@@ 3.2 Prompt and scoring modifications
LLMs exhibit high sensitivity to variations in prompt formatting Sanh et al. (2021); Mishra et al. (2022), forcing benchmark developers to unify prompt templates within the same evaluation scheme. However, it remains unknown if certain models have an affinity towards any specific prompt templating style. It is unclear how benchmarking prompt choices advantage/disadvantage different models. In addition to that, scoring style may change depending on how we are prompting the context of a query. We distinguish three major categories of scoring methods for MCQs.
@@
Cloze scoring: Prompt templates are structured as question followed by a single answer choice. Maximum normalized likelihood scores over all answer choices defines the prediction. Used in Clark et al. (2018a).
@@
To assess the impact of subtle token alterations in prompt instructions, we conduct experiments on (i) removing question subject information, and (ii) adding "Correct" alongside the answer. These targeted changes aim to identify the robustness in response to certain tokens, particularly when they carry crucial information, as well as to evaluate the influence of contextual bias introduced by minor modifications of the instruction text.
@@ 3.3 In-context knowledge manipulation
Under this setting, we attempt to measure model and benchmark robustness in the few-shot setting by testing the entire spectrum of knowledge injected in the few-shot examples. In particular, we observe how performance changes under trivial settings where the correct answer is provided in the context, as well as when the examples are irrelevant to the question.
@@
This setting is the opposite, the target question is provided with an incorrect answer as an example. It is challenging as the model must ignore the context and determine the correct answer independently.
@@
We replace few-shot examples with simple questions the model is known to be able to answer (typically related to the language/text of the question itself). The only information conveyed by the examples is related to formatting Soltan et al. (2022). We create three versions of these questions and answers using GPT-4 and ensure the model can answer them correctly (as shown in Figure A.1).
@@
Instead of providing examples from the same subject as the target question, we add out of domain questions (from another subject) as the few-shot examples. This setting corresponds to a difficulty level between the original format and providing trivial examples.
@@ 4 Experiments
MMLU Hendrycks et al. (2020) is a commonly used benchmark for comparing LLMs, consisting of 57 subjects spanning four domains: humanities, STEM, social sciences, and others. Each subject includes at least 100 multiple choice questions with 4 answer choices. The entire benchmark contains 14,042 questions.
@@
Unless otherwise stated, the reported score for each experiment/model combination on MMLU is the mean accuracy across all 14,042 questions. All tested model tokenizers encode the multiple choice answers as single tokens. Hence, the accuracy is equivalent to the normalized accuracy. All baseline and modified MMLU benchmarks were performed using the LM Evaluation Harness Gao et al. (2023) library. Their implementation of MMLU measures the log-likelihood of each of the answer tokens [’A’, ’B’, ’C’, ’D’] after the input prompt and chooses the letter with the highest probability as the model’s answer.
Some of our experiments require permuting the answer choice order, however, this can be confusing for questions where the answer choices are dependent on their position, such as “D. All of the above.”, or reference other choices, such as “C. Both A and B.”. To circumvent this dependency, we manually inspected and modified the questions from three subjects to ensure their answers are permutation independent for a subset of our experiments. The modified subjects are: college chemistry, college mathematics, and global facts.
@@ 5.1 MCQ benchmarks are not robust to perturbations
As shown in Figure 1, there exist perturbations which cause dramatic shifts in the order of models with respect to commonly accepted leaderboard rankings. We find a significant number of small perturbations demonstrate this effect, while other perturbations are more benign.
@@
Shuffling/changing the presented order of the choices, swapping choice symbols, and alternative scoring methods all cause major shifts to the rankings (determined by thresholding ). For example, in a controlled experiment where we randomly shuffle the answer choices presented to the models (Table 1). Five out of 11 models change in ranking after the perturbation and drops to 0.564. A similar pattern is seen for perturbations like fixing the correct answer to a particular position (Table 2), replacing the default choice symbols with other sets (Figure 4), and alternative scoring methods (Figure 7).
Some models elicit this behavior much more strongly. For example, we observe that Yi-6b drops from third place to seventh or eight place under some benchmark perturbations in the group of 11 models we tested. Other models in the same size range are more stable (e.g., Mistral-7b, Llama2-7b), not shifting more than one or two ranks under all perturbations. The reasons for this are not clear, but could indicate overfitting to aspects of the benchmark style. Since training data for these models is not public, it is difficult for us to verify this hypothesis.
@@ 5.3 Another source of bias: scoring bias
Beyond the ordering of choices and the symbols associated with them, LLMs exhibit varying amounts of bias under the choice of scoring method for MCQs. We studied the three scoring methods described in Section 3.2: symbol scoring, cloze scoring, and hybrid scoring. Symbol scoring has become the dominant method for evaluating LLMs on MCQs, largely due to the high accuracy achieved by LLMs Robinson et al. (2023). This, however, comes at the cost of high selection bias. Cloze scoring can essentially eliminate bias since the choices are never presented to the model but LLMs tend to score poorly using this method. This also does not reflect a true MCQ setting. Figure 7 and Table A.13 detail the results of these experiments..
@@ 5.4 Minor few-shot and prompt changes have little effect on benchmark rankings
We ran several experiments to assess the effect of knowledge provided in-context on model performance and rankings. We find that changing the informativeness of in-context examples, e.g. providing irrelevant/trivial examples (Tables A.17-A.19) or examples from subjects other than the target subject (Figure A.4, Tables A.24- A.25), slightly changes performance across models and reduces bias compared to zero-shot settings but does not change rankings drastically. This finding leads us to conclude that adding few shot examples to benchmark evaluations can help reduce, but not eliminate, leaderboard sensitivity.

### v2

@@ Abstract
Large Language Model (LLM) leaderboards based on benchmark rankings are regularly used to guide practitioners in model selection. Often, the published leaderboard rankings are taken at face value — we show this is a (potentially costly) mistake. Under existing leaderboards, the relative performance of LLMs is highly sensitive to (often minute) details. We show that for popular multiple-choice question benchmarks (e.g., MMLU), minor perturbations to the benchmark, such as changing the order of choices or the method of answer selection, result in changes in rankings up to 8 positions. We explain this phenomenon by conducting systematic experiments over three broad categories of benchmark perturbations and identifying the sources of this behavior. Our analysis results in several best-practice recommendations, including the advantage of a hybrid scoring method for answer selection. Our study highlights the dangers of relying on simple benchmark evaluations and charts the path for more robust evaluation schemes on the existing benchmarks. The code for this paper is available at https://github.com/National-Center-for-AI-Saudi-Arabia/lm-evaluation-harness.
@@
Norah A. Alzahrani††thanks: Core contributor, Hisham Abdullah Alyahya††footnotemark: , Yazeed Alnumay, Sultan Alrashed, Shaykhah Z. Alsubaie, Yousef Almushayqih, Faisal Abdulrahman Mirza, Nouf M. Alotaibi Nora Al-Twairesh, Areeb Alowisheq, M Saiful Bari, Haidar Khan††footnotemark: ††thanks: Corresponding author: haidark@sdaia.gov.sa National Center for AI (NCAI), Saudi Data and AI Authority (SDAIA) Riyadh, Saudi Arabia
@@ 1 Introduction
The advent of transformer-based Large Language Models (LLMs) (OpenAI, 2023; Deepmind, 2023; Anthropic, 2023; Anil et al., 2023; Touvron et al., 2023) has led to a generational leap in generative models, enabling interaction with computing devices through natural language. This advancement encompasses improvements that have rendered many earlier benchmarks and leaderboards obsolete (Laskar et al., 2023; Shen et al., 2023), leading to the compilation of more challenging and comprehensive tests. However, the current generation of leaderboards still does not satisfy many of the requirements of researchers and practitioners looking to build on LLMs Ethayarajh and Jurafsky (2021); Dehghani et al. (2021). Since LLMs are extremely expensive to both train and inference, selecting the LLM (or LLM training recipe) is often the most costly decision for the entire project. Stable leaderboards are critical to making the right decision.
@@
Leaderboards based on multiple choice questions (MCQ) for evaluation (Wang et al., 2018, 2019; Nie et al., 2019; Zhong et al., 2023; Hendrycks et al., 2020) present both convenience and significant limitations (Pezeshkpour and Hruschka, 2023; Zheng et al., 2023). While MCQs offer an automated and quantifiable means to assess certain aspects of model ability (e.g., knowledge), they fall short as a stable means to measure performance. Figure 1 demonstrates the instability of the leaderboard ranking of one popular benchmark, Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020), under small perturbations.
@@
Moreover, the reliance on MCQs raises concerns about the models being overfit to these benchmarks, potentially excelling in structured tests while lacking real-world applicability. This discrepancy highlights the need for more holistic and diverse evaluation methods that transcend the simplicity of MCQs Liang et al. (2023). It also prompts critical reflection on how these models might inadvertently be trained to achieve high scores through spurious correlations, pattern recognition, and optimization for specific question formats rather than genuine language comprehension or knowledge. As LLMs continue to evolve, it is imperative to develop evaluation frameworks that can more accurately assess their abilities in a way that mirrors the complexity of real-world use.
@@
Despite being widely used, benchmarking with MCQs has turned out to be anything but simple. It requires the full synchronization of evaluation frameworks and results often vary wildly due to nuanced differences. For example, minor changes in prompting and scoring can produce invalid results for particular LLMs111https://huggingface.co/blog/evaluating-mmlu-leaderboard. Recent studies have investigated the issue of sensitivity in evaluating LLMs, with some demonstrating that LLMs are susceptible to the ordering of answer choices and bias towards specific tokens/symbols (Zheng et al., 2023; Pezeshkpour and Hruschka, 2023; Lu et al., 2022), and others exploring different prompt modifications and their effects in benchmarking LLMs (Mizrahi et al., 2024; Sclar et al., 2023; Weber et al., 2023; Zhao et al., 2021).
In this work, we conduct a broad range of minor perturbation experiments to MCQ benchmarks and observe the disruption it causes to model rankings on leaderboards. We also take additional steps to precisely identify the limitations of LLMs on this measurement approach.
@@
We demonstrate that some categories of modifications do not affect the benchmark rankings.
@@ 2 LLM Evaluation with MCQs
Mechanically, testing LLMs with MCQs is accomplished by presenting the question along with the answer choices to the model and selecting the choice deemed most probable by the model. Although this setup appears straightforward, LLMs react in unpredictable ways to formatting and other minor changes to the questions or the answers. LLM performance on an MCQ test can change with the introduction of an extra space (e.g., between the question and answer) or adding an additional instructional phrase (e.g., "Choices:"). In addition to this brittleness, Pezeshkpour and Hruschka (2023) found changes to the order in which answer choices are presented to GPT4 and instructGPT can change the model’s prediction.
@@
Our main aim is to quantify how these small perturbations/variations change the rankings of a set of models on a particular benchmark. As MCQ benchmarks-based leaderboards are often used to compare models and guide model selection, we investigate the robustness of benchmarks for this purpose. Figure 1 demonstrates how existing benchmarks exhibit significant undesirable shifts in rankings under small perturbations.
@@ 3 Methods
In this section, we describe and justify the perturbations we apply in each category. We note that some MCQ test changes, like modifying the order of answer choices, can change performance even for humans but the effect is typically not pronounced (Lions et al., 2021). In general, our modifications are designed to be small perturbations to the MCQ and prompts that should not affect performance. The exception to this is some of the in-context knowledge manipulations described in Section 3.3, which are designed to improve or degrade performance drastically.
@@ 3.1 Answer choice format and ordering
In light of earlier findings related to selection bias Zheng et al. (2023); Pezeshkpour and Hruschka (2023); Lu et al. (2022), we investigate the effects of changes to the presented order of answer choices and changes to the symbols associated with answer choices.
@@
Our first study aims to uncover how dependent MCQ benchmark performance and rankings are on the original ordering of the answer choices. We apply two simple schemes to randomly change the order of answer choices presented to the model: (i) swapping choices using a fixed set of swaps for all questions and (ii) randomly assigning new positions to each choice while ensuring each choice is moved to a different position.
@@
In this setting, the correct answer choice is set to a fixed position across the entire test to measure bias toward predicting answers at particular positions. For zero-shot, we simply set the correct answer choice to each of the positions in turn.
In the few-shot case, we examine the influence of biasing the correct answers in the examples to the model’s inherent bias to particular positions. For each question, we fix the correct answer of the examples to each position in turn. We then modify the test question in two ways: (i) unchanged answer choices and (ii) correct choice fixed to the same position as examples. This setup is shown in Figure 2.
@@
The symbols used for the answer choices (e.g. A, B, C, D) also play a role in model bias Zheng et al. (2023). Thus, we experiment with replacing the symbols with alternative and less common tokens. This aims to decouple the bias to particular positions from the bias to symbols or the relative ordering in natural symbols. We replace [’A’, ’B’,’ C’, ’D’] with the following two sets of symbols:
@@ 3.2 Prompt and scoring modifications
LLMs exhibit high sensitivity to variations in prompt formatting Sanh et al. (2021); Mishra et al. (2022), forcing benchmark developers to unify prompt templates within the same evaluation scheme. However, it remains unknown if certain models have an affinity towards any specific prompt templating style. It is unclear how benchmarking prompt choices advantage/disadvantage different models. In addition to that, the scoring style may change depending on how we are prompting the context of a query. We distinguish three major categories of scoring methods for MCQs.
@@
Cloze scoring: Prompt templates are structured as a question followed by a single answer choice. Maximum normalized likelihood scores over all answer choices define the prediction. Used in Clark et al. (2018a).
@@
To assess the impact of subtle token alterations in prompt instructions, we conduct experiments on (i) removing question subject information and (ii) adding "Correct" alongside the answer. These targeted changes aim to identify the robustness in response to certain tokens, particularly when they carry crucial information, as well as to evaluate the influence of contextual bias introduced by minor modifications of the instruction text.
@@ 3.3 In-context knowledge manipulation
In this category, we attempt to measure model and benchmark robustness in the few-shot setting by testing the entire spectrum of knowledge injected in the few-shot examples. In particular, we experiment under the following settings:
@@
This setting is the opposite of the former. The target question is provided with an incorrect answer as an example. It is challenging as the model must ignore the context and determine the correct answer independently.
@@
We replace few-shot examples with simple questions the model is known to be able to answer (typically related to the language/text of the question itself). The only information the examples convey is related to formatting Soltan et al. (2022). We create three versions of these questions and answers using GPT-4 and ensure the model can answer them correctly (as shown in Figure 8).
@@
Instead of providing examples from the same subject as the target question, we add out-of-domain questions (from another subject) as the few-shot examples. This setting corresponds to a difficulty level between the original format and providing trivial examples.
@@ 4 Experiments
MMLU Hendrycks et al. (2020) is a commonly used benchmark for comparing LLMs, consisting of 57 subjects spanning four domains: humanities, STEM, social sciences, and others. Each subject includes at least 100 multiple-choice questions with 4 answer choices. The entire benchmark contains 14,042 questions (Tables A.29 and A.30 have a breakdown of the MMLU subjects and their distributions).
@@
Unless otherwise stated, the reported score for each experiment/model combination on MMLU is the mean accuracy across all 14,042 questions. All tested model tokenizers encode the multiple-choice answers as single tokens. Hence, the accuracy is equivalent to the normalized accuracy. All baseline and modified MMLU benchmarks were performed using the LM Evaluation Harness Gao et al. (2023) library. Their implementation of MMLU measures the log-likelihood of each of the answer tokens [’A’, ’B’, ’C’, ’D’] after the input prompt and chooses the letter with the highest probability as the model’s answer.
Some of our experiments require permuting the answer choice order (Table 1, A.8, A.9, and A.10), however, this can be confusing for questions where the answer choices are dependent on their position, such as “D. All of the above.”, or reference other choices, such as “C. Both A and B.”. To circumvent this dependency, we manually inspected and modified the questions from three subjects to ensure their answers are permutation-independent for a subset of our experiments. The modified subjects are college chemistry, college mathematics, and global facts.
@@ 5.1 MCQ benchmarks are not robust to perturbations
As shown in Figure 1, there exist perturbations that cause dramatic shifts in the order of models with respect to commonly accepted leaderboard rankings. We find a significant number of small perturbations demonstrate this effect, while other perturbations are more benign.
@@
Shuffling/changing the presented order of the choices, swapping choice symbols, and alternative scoring methods all cause major shifts to the rankings (determined by thresholding ). For example, in a controlled experiment using a subset of MMLU, we randomly shuffle the answer choices presented to the models (Table 1). 5 out of 11 models change in ranking after the perturbation and drops to 0.564. A similar pattern is seen for perturbations like fixing the correct answer to a particular position (Table 2), replacing the default choice symbols with other sets (Figure 4), and alternative scoring methods (Figure 7).
Some models elicit this behavior much more strongly. For example, we observe that Yi-6b drops from 3rd place to 7th or 8th place under some benchmark perturbations in the group of 11 models we tested (namely, the rare symbol and cloze perturbations). Other models in the same size range are more stable (e.g., Mistral-7b, Llama2-7b), not shifting more than one or two ranks under all perturbations. The reasons for this are unclear but could indicate overfitting to aspects of the benchmark style. Since training data for these models is not public, it is difficult for us to verify this hypothesis.
@@ 5.3 Another source of bias: scoring bias
Beyond the ordering of choices and the symbols associated with them, LLMs exhibit varying amounts of bias under the choice of scoring method for MCQs. We studied the three scoring methods described in Section 3.2: symbol scoring, cloze scoring, and hybrid scoring. Symbol scoring has become the dominant method for evaluating LLMs on MCQs, largely due to the high accuracy achieved by LLMs Robinson et al. (2023). This, however, comes at the cost of high selection bias. Cloze scoring can essentially eliminate bias since the choices are never presented to the model, but LLMs tend to score poorly when using this method. This also does not reflect a true MCQ setting. Figure 7 and Table A.13 detail the results of these experiments.
@@ 5.4 Minor few-shot and prompt changes have little effect on benchmark rankings
We ran several experiments to assess the effect of the initial prompt on model performance and rankings. We find that changing the informativeness of in-context examples, e.g. providing irrelevant/trivial examples (Figure 8, Tables A.17-A.19) or examples from subjects other than the target subject (Figure 9, Tables A.24- A.25), slightly changes performance across models and reduces bias compared to zero-shot settings but does not change rankings drastically. This finding leads us to conclude that adding few shot examples to benchmark evaluations can help reduce, but not eliminate, leaderboard sensitivity.
We also experiment with removing subject information from instructions and ending the prompt with "Correct Answer:" instead of "Answer:" (Figure 10, Tables A.20-A.23). We see little changes () in these prompt modification experiments.